OpenStudy (anonymous):

Find the sum of the measures of the interior angles of a convex 30-gon.

OpenStudy (cwrw238):

there is a formula to find this directly but i can never remember it for some reason. I do these by using the fact that the sum of the exterior angles of all polygons add up to 360 degrres.

OpenStudy (cwrw238):

so for a convex regular 30 sided polygon each exterior angle = 360/30 = 12 degrees so each interior angle = 180-12 = 168 degrees

OpenStudy (alexandervonhumboldt2):

180*(30-2)

OpenStudy (alexandervonhumboldt2):

180*28

OpenStudy (alexandervonhumboldt2):

now count

OpenStudy (cwrw238):

so multiply this by 30 and you have your answer

OpenStudy (alexandervonhumboldt2):

180*(n-2) wjere n is number of sides

OpenStudy (alexandervonhumboldt2):

wait nononono

OpenStudy (alexandervonhumboldt2):

imagine this is less that sumof angle of triangle@@!!!!

OpenStudy (cwrw238):

yes - both methods give the same answer

OpenStudy (alexandervonhumboldt2):

180*(n-2) I know this!!!!!!! where n is now in this situation is 30

OpenStudy (cwrw238):

why do the posts sometimes come out in the wrong order!!???

OpenStudy (cwrw238):

my last 2 did just that!!

OpenStudy (alexandervonhumboldt2):

i dont know yeah i noticed that tool ooks like we are crazy

OpenStudy (anonymous):

The posts are a little mixed up. Is 168 the final answer?